1)Does y vary directly, inversely or neither with x? If so, write the equation for the variation. (It's a chart but I'll type them for x and y)
x - -2, 3, 8
y- -2.8, 4.2, 11.2
2)Does y vary directly, inversely or neither with x? If so, write the equation for the variation.
x- 4, 5, 11
y- 3, 4.5, 13.5.

Answer by Theo(13342)

if they vary directly, then k = y/x
if they very inversely, then k = y*x

this is because the equation for direct variation is y = k*x
solve for k and you get k = y/x

the equation for inverse variation is y = k/x
solve for k and you get k = y*x

k has to be the same for all of the pairs, not just some of them.

your first series is:

x = -2, 3, 8
y = -2.8, 4.2, 11.2

we'll look for direct variation first.

-2.8 / -2 = 1.4
4.2 / 3 = 1.4
11.2 / 8 = 1.4

it's a direct variation equation.

your second series is:

x = 4, 5, 11
y = 3, 4.5, 13.5.

we'll look for direct variation first.

3/4 = .75
4.5 / 5 = .9

it's not a direct variation.

let's look for inverse.

3 * 4 = 12
4.5 * 5 = 22.5

it's not an inverse variation.

this one is neither.